### v2

# Investigating Modality Contribution in Audio LLMs for Music

Abstract
Audio Large Language Models (Audio LLMs) enable human-like conversation about music, yet it is unclear if they are truly listening to the audio or just using textual reasoning, as recent benchmarks suggest. This paper investigates this issue by quantifying the contribution of each modality to a model’s output. We adapt the MM-SHAP framework, a performance-agnostic score based on Shapley values that quantifies the relative contribution of each modality to a model’s prediction. We evaluate two models on the MuChoMusic benchmark and find that the model with higher accuracy relies more on text to answer questions, but further inspection shows that even if the overall audio contribution is low, models can successfully localize key sound events, suggesting that audio is not entirely ignored. Our study is the first application of MM-SHAP to Audio LLMs and we hope it will serve as a foundational step for future research in explainable AI and audio.
Index Terms—  Audio LLMs, Explainable AI, Shapley Values, MM-SHAP, Modality Contribution
1 Introduction
Audio Large Language Models (Audio LLMs) aim to expand the capabilities of LLMs by incorporating audio information into their reasoning [5]. While the number of proposed models continues to grow [3, 13, 4, 8, 26] difficulty lies in assessing free-form text outputs, due to the unstructured nature of the predictions [12]. Additionally, evaluation benchmarks must go beyond measuring natural language understanding to explicitly test a model’s ability to reason using audio. Multiple evaluation datasets have been proposed that assess model performance in tasks such as captioning [1] and compositional reasoning [7]. The lack of consistent evaluation methods inspired researchers to build new comprehensive benchmarks to evaluate and compare Audio LLMs, such as MuChoMusic [27] and the Massive Multi-Task Audio Understanding and Reasoning Benchmark (MMAU) [23]. These datasets feature multiple-choice questions, with the goal that the model will choose the most appropriate answer based on the input audio. The main advantages of this evaluation method are scalability and simplicity, since the model’s answers are matched against predefined options rather than parsed from open-ended text. Furthermore, the questions are designed to test various dimensions of music understanding, such as historical knowledge, music structure, and rhythm.
A surprising outcome is that audio-text models perform poorly on these benchmarks, despite achieving strong results on tasks such as audio captioning and description [27, 23], although the reasons for this remain unclear. Experiments that replace the input audio with random noise or silence result in negligible changes to model performance [27], suggesting that the models’ use of the audio modality may be limited. While these outcomes indicate a potential imbalance in the modalities (i.e., the text modality is the dominant contributor to the model’s output), the benchmarks primarily rely on indirect metrics, like accuracy, which can be misleading. Accuracy-based evaluations fail to capture cases where a model uses relevant information from a modality but still produces an incorrect prediction.
Similar modality-imbalance issues have been reported in Vision LLMs, where unimodal models can often achieve comparable accuracy to a multimodal one on a vision-language task [6], suggesting unimodal collapse. To better understand this phenomenon without relying on performance metrics, [19, 20] proposed MultiModal-SHAP (MM-SHAP). They utilize Shapley values [24], a post-hoc explainability technique to quantify the relative contribution of each modality in a model’s prediction, irrespective of its accuracy. We investigate MM-SHAP’s adaptability to the modality of audio. Particularly, our main contributions are that we:
  1. 1.
    Adapt MM-SHAP [19] to inspect how Audio LLMs are using each modality in different tasks;
  2. 2.
    Investigate two well-known Audio LLMs (Qwen-Audio [3] and MU-LLaMA [13]) in multiple-choice Q&A using the open-source MuChoMusic benchmark dataset [27]; and
  3. 3.
    Examine how the Audio LLMs use the two modalities. We show that the usage of text is higher for multiple-choice questions, aligning with results from Vision LLMs. We also demonstrate that good performance on MuChoMusic does not imply balanced modality contributions and vice versa.
Our aim is to gain insight into how much these models employ each modality to reason over music, and how the task and the MM-SHAP design choices affect these interactions.
2 Related Work
Fig. 1: We compute Shapley values by masking all combinations of inputs (approximated via random permutation), and averaging the change in the logits of the base answer (unmasked inference indicated in solid line). We mask text tokens and audio waveform segments.
Within explainability techniques in machine learning, there is a family of “post-hoc” methods whose aim is to analyze how each input feature contributes to a given model output via input perturbations. These include approaches like LIME (Local Interpretable Model-agnostic Explanations) [22], that explain a prediction by approximating it locally with an interpretable model (e.g., linear regression or a decision tree). To the best of our knowledge, most of the post-hoc techniques applied to music use LIME or LIME-variations [9],[17]. Recently, [25] expanded LIME methods to investigate music understanding models by analyzing the most important features for genre classification from both text and audio modalities. As for modality contribution, the DIsentangled Multimodal Explanations (DIME) [15] was proposed as a visualization technique to gain insights into each modality’s contribution to the model’s output. DIME disentangles a model into unimodal and multimodal contributions and then produces visualizations on each sub-model using LIME. Because of its focus on visual explanations, DIME is hard to use as a quantitative measure of modality contribution.
A common strategy for assessing modality contribution is to conduct ablation studies by removing or replacing one of the modalities with a random input. For example, [27] swaps the audio with silence, white noise, or another audio and evaluates how much the accuracy changes. The main limitation of this approach is that completely removing a modality obscures how the model uses each modality and their interactions, and to what extent its output depends on the missing modality. As argued in [10], comparing the unimodal and multimodal performances does not explain how the model uses cross-modality interactions.
To address the limitations of ablation studies and provide stronger theoretical guarantees than surrogate models like LIME, another family of “post-hoc” methods based on Shapley values has gained prominence. Shapley values are the only feature attribution method that satisfies theoretical guarantees [18]. To the best of our knowledge, the only exploration of Shapley values and music audio was performed in [2]. Recently, Parcalabescu and Frank proposed MM-SHAP [19], which quantifies modality contribution in text-vision models by adapting Shapley values to measure how input perturbations across modalities affect the model’s output.
3 Method
3.1 Shapley Values and Feature Contribution
Shapley values were first proposed in the context of game theory to estimate how much each player contributes to the overall outcome of a cooperative game [24]. They were adapted for explaining machine learning methods by [14] in the SHAP (SHapley Additive exPlanations) framework, where features act as players contributing to a model’s output. The Shapley value $\phi$ for a feature $j$ and token $t$ is given by
$\phi_{j,t}=\sum_{S\subseteq\{1,\cdots,n\}/\{j\}}\frac{f_{t}(S\cup\{j\})-f_{t}(S)}{\gamma},$ (1)
where $S$ represents subsets of features excluding the feature $j$, $f_{t}(S)$ is the model’s output for the subset $S$ and token $t$, $n$ defines the total number of features. Therefore, $\phi_{j,t}$ represents how much feature $j$ contributes to the output token $t$. The normalizing factor $\gamma=\frac{n!}{|S|!(n-|S|)!}$ weights the marginal contribution of each feature fairly across all possible subsets of features. A subset’s marginal contribution is weighted proportionally to how often that subset appears when averaging over all permutations. Shapley values satisfy the theorical guarantees of symmetry, efficiency, dummy, and additivity, which together constitute a fair payout [18].
The exact computation of Shapley values requires evaluating all possible subsets of features. However, as the number of features grows, the computation of Equation 1 has an exponential time complexity of $O(2^{n})$. This means that, for machine learning models with a large numbers of features, the exact computation of Shapley values is too expensive. In practice, Shapley values are approximated by sampling feature subsets, e.g., using KernelSHAP or PermutationSHAP from the SHAP framework [14]. In this work, we use PermutationSHAP, which approximates Shapley values by averaging each feature’s marginal contribution across random permutations, measuring the change in the model’s output when the feature is added to preceding subsets. When estimated using $m$ randomly sampled subsets, PermutationSHAP has a complexity of O(mn)O(m*n). We use the default value of $m=10$ from the shap library.
3.2 Features and Masking
For audio we mask the raw waveform by zeroing out short waveform segments. The advantage of masking raw inputs rather than intermediate audio representations is that it prevents information leakage (e.g., in case the audio tokens are computed with temporal overlap or redundancy), and allows for further interpretability of how the masked parts of the input affect the output.
Our audio masking adapts dynamically to text length, aiming for a balanced number of features from each modality. This is accomplished by adjusting the window size such that the quotient of dividing the audio length by the number of windows equals the number of text tokens ($n_{A}=n_{T}$), following [19]. For example, in a 10-second audio clip and 100 text tokens, the masked windows are approximately 100ms long.
When masking text tokens, we replace them with the [MASK] token. We do not mask certain special tokens such as “audio indicators” which signal that audio features follow (<audio>, <audio_padding>, #Audio), and question and answer indicators (<|question|>, <|answer|>).
3.3 Computing Model Scores
For a classification task such as the ones explored in [19], the model’s score for feature subset $S$ given a class $c$ is defined as the output probability of the model $f_{c}(S)$ for that class. The Shapley values are then computed by measuring how the probabilities change as the input features are masked. Scores can be computed at the class-level or averaged across classes for a “global” effect. For a generation task, this approach would be too costly as the “classes” involved span the whole token vocabulary of the LLM. Instead, we leverage the ideas in [21] to develop the methodology illustrated in Figure 1. First, we obtain a set of tokens $T$ corresponding to the model’s answer without masking any modality (i.e., “(B) Doorbell” in Figure 1). We save the indices of the tokens of the generated answer, and then compute $f_{t}(S)$ by summing the changes in the logits for each token $t\in T$ after masking parts of the input (in both text and audio). This way we measure the changes of the model with respect to the baseline, unmasked answer.
Fig. 2: QwenAudio MC-NPI results for MusicCaps track QK-mjNg8cPo: the output is “The sound effect that can be heard in the piece is a bell sound effect”. The image shows the modality contribution for output token “bell”, i.e., $\Phi_{A,t}$ and $\Phi_{T,t}$, where $t=\text{bell}$. The top section highlights the most important text tokens from the input question (black = highest contribution, with a threshold of 80% of the maximum Shapley value for readability). The bottom section shows the audio waveform with its corresponding Shapley value contributions: absolute value, positive, and negative components. Darker colors mean higher contribution to the model’s output token.
3.4 Modality Contribution
While Equation 1 captures the impact of a single feature $j$, we are interested in understanding how sets of features, grouped by modality, contribute to the model’s output. Following the proposal of [19], we define each modality contribution as
$\Phi_{A}=\sum_{t}^{T}\sum_{j=1}^{n_{A}}|\phi_{j,t}|$ $\Phi_{T}=\sum_{t}^{T}\sum_{j=n_{A}+1}^{n_{A}+n_{T}}|\phi_{j,t}|,$ (2)
where $\Phi_{A}$ and $\Phi_{T}$ represent the scores of all audio and text features, and $n_{A}$ and $n_{T}$ are the number of audio and text tokens, respectively. We note here that audio and text features are concatenated. Shapley values measure positive and negative contribution amounts. However, for measuring modality contribution, we are interested in how much a feature contributed to the expected output. Therefore, we use the absolute value of the Shapley value $\phi_{j,t}$. To directly compare modality influence, we define T-SHAP (Textual SHAP) and A-SHAP (Audio SHAP) as the normalized ratios of each modality’s total contribution. For a perfectly balanced model, both T-SHAP and A-SHAP would equal $0.5$.
$\text{A-SHAP}=\frac{\Phi_{A}}{\Phi_{T}+\Phi_{A}}$ $\text{T-SHAP}=1-\text{A-SHAP}.$ (3)
4 Experiments
Data Our experiments focus on the MuChoMusic benchmark [27], as we have access to both the audio and the questions’ answers. MuChoMusic is made of 1,187 multiple-choice questions validated by human annotators and associated with 644 music tracks sourced from the SongDescriberDataset (SDD) and MusicCaps. We limit our experiments to MusicCaps tracks to make a fair comparison between models, since Qwen-Audio focuses on the first 30s of the audio [^1] https://github.com/QwenLM/Qwen-Audio?tab=readme-ov-file#quickstart, while MU-LLaMA processes entire tracks. MusicCaps questions constitute approximately 71% of MuChoMusic.
Models Our experiments compare two Audio LLMs, which represent the best and second-worst performing models on MuChoMusic. MU-LLaMA [13] uses MERT-v1-330M as the audio encoder and LLaMA 2 7B as the language encoder, then trains an adapter to map the MERT output to the language model. All audio is resampled to 24 kHz before encoding. The model is trained with the MusicQA dataset, which is compiled from the MusicCaps [1] and MagnaTagATune [11] datasets. Qwen-Audio [3] uses Whisper-large-v2 as its base audio encoder and Qwen-7B as its base LLM. Audio is resampled to 16 kHz prior to encoding. We use Qwen-Audio-Chat for our experiments as it is fine-tuned to answer questions.
Experiments To explore the effect of the number of text tokens on the audio usage, we investigate the models in two different setups:
  1. 1.
    Multiple-Choice with Previous Instructions (MC-PI): We use the original input of MuChoMusic containing an in-context example followed by the actual multiple-choice question we want the model to answer.
  2. 2.
    Multiple-Choice with No Previous Instructions (MC-NPI): We omit the in-context example (i.e., fewer text tokens).
For each task and model, we report A-SHAP scores and accuracy, calculated with the code provided in the MuChoMusic GitHub repository. We kept all system instructions with their default values (“You’re a reliable assistant, follow these instructions.”). Unlike [19], we did not mask system instructions, only the question itself.
We also performed a qualitative analysis to investigate how the Shapley values behave for questions requiring single sound event vs. those requiring longer-term audio understanding to generate the answer. We looked at each model and each experiment, filtered only questions with correct answers and selected 5 samples with the highest A-SHAP values, and 5 with the lowest. We removed duplicates, a sample with corrupted audio, and a question with no correct distractor, resulting in 35 unique Q&A. From this set, we first identified questions answered by a single sound event. For these, we manually annotated the relevant audio segment, as in Figure 2. We observed that most of these questions still required a long-term understanding of the audio. To find more single-sound-event examples, we searched for keywords that were common in such questions (e.g., “interrupted”, ”at the beginning“, ”sound at the end“). We identified a total of 13 Q&A pairs. All of our experiments and code are available on our GitHub [^2] https://github.com/giovana-morais/2025_investigating_mmshap.
5 Results and Discussion
Table 1: Accuracy and average A-SHAP for the two experiments.
Accuracy A-SHAP
MC-PI MC-NPI MC-PI MC-NPI
MU-LLaMA $0.30$ $0.32$ $0.50\pm 0.02$ $0.47\pm 0.02$
QwenAudio $0.44$ $0.47$ $0.23\pm 0.02$ $0.21\pm 0.02$
Table 1 shows that the more accurate model, Qwen-Audio, relies less on audio, whereas MU-LLaMA uses both modalities in a balanced manner. Indicating that there is no relationship between accuracy and modality usage. This was further reinforced by analyzing A-SHAP values grouped by correct and incorrect questions, which showed no trend. Examining questions broken down by knowledge dimension according to MuChoMusic categories (e.g. knowledge, reasoning) also showed no significant effect in the A-SHAP value. We think that this lack of difference is consistent with the findings in [28], where the authors show that LLMs perform as good on MuChoMusic as Audio LLMs, indicating that the audio modality is not leveraged as expected in the answering process. They show that this is caused by MuChoMusic approach to designing the multiple-choice distractors which, as a result, any LLM capable of eliminating the two unrelated answers can achieve an accuracy of 50% by randomly selecting the answer from the two remaining options. What the results of [28] suggest is that what MuChoMusic is, in fact, testing the reasoning skills of an LLM model instead of the perceptual capabilities of Audio LLMs. To verify whether the low audio usage is extended to tasks other than multiple-choice, we additionally evaluated the best-performing model (Qwen-Audio) by asking it to describe the audio clips, with the same number of audio tokens as MC-PI, for fairness.The average A-SHAP increased to 0.73%, suggesting stronger audio usage and significant task influence.
As we have seen little change in the audio usage between correct and incorrect examples, we decided to further explore the positive and negative contributions ($\phi_{j,t}$), together with the absolute value that is in fact used to compute MM-SHAP. We analyze the songs as described in Section 4. From the qualitative analysis, we found that in Qwen-Audio the audio features have much lower magnitude than text features, which aligns with the low audio usage. But this does not mean that the audio is being discarded. For each specific token $t\in T$, we see that the positive contributions often align with the audio region that contains the ground truth, and also with the input token that represents that event, as shown in Figure 2 for the token “bell”. We note that the same behavior was found in the other questions with single-sounding events. It is also possible to see other strong activations, and those are not related to the correct answer (in Figure 2, they represent a piano note). We hypothesize that in such cases, the information is useful for the model’s internal mechanisms, but not necessarily are plausible to the data explanation. In examples that require longer-term explanations, the activations are spread throughout the audio. In MU-LLaMA, the audio and text features had similar magnitude overall, but we noticed that the audio features with high activations were spread throughout the audio in long-term and single-sounding questions. Although there were positive contributions in the correct regions, they were not as concentrated as in Qwen-Audio. Also, MU-LLaMA activations showed negative contributions as strong as the positive, which affect the final A-SHAP value that takes into account only the magnitude of the feature importance. We provide additional interactive plots comparing all models and experiments in the project’s demo page.
To the best of our knowledge, this is the first time modality contribution is measured in Audio LLMs, as inspected through Shapley values. In this work, we followed the original proposal by [19] closely. However, we discuss below the factors that, in our opinion, limit the application of MM-SHAP for Audio LLMs as is, and venues for improvement and future research.
A primary challenge is the lack of ground truth for qualitative analysis. Our analysis was restricted to questions with easily localizable sound events; annotating questions about melody, instrumentation, or structure is harder, yet those question types are the most frequent when evaluating music understanding systems, which makes reliable annotation an open question by itself. Furthermore, the audio segment size, which is tied to the number of text tokens, may be too small to contain perceptually meaningful information, complicating our cross-experiment comparison. For the audio masking, we zeroed the audio segments, but other alternatives include replacing them with white noise or an average between neighbor segments, as in [2] and [16]. Determining an optimal audio window size and the optimal audio mask are open research problems, especially if accounting for temporal and spectral information. This is especially important because the chosen mask must be a “neutral” representation of the data. While MM-SHAP and Shapley values can help us understand how much audio is used in current models, the tool is not enough to hint us about how the audio is used internally in the model, as is the case with post-hoc explanation methods.
6 Conclusion
In this work, we adapt MM-SHAP to measure modality contributions in Audio LLMs, analyzing how audio and text are used in perception tasks. Through a systematic evaluation of Qwen-Audio and MU-LLaMA, we discovered that the best-performing model, Qwen-Audio, relied significantly more on its text modality. While text processing appears to drive overall performance, our qualitative analysis confirms that the audio modality is not discarded, suggesting a gap between listening to the audio and reasoning about it. Our findings also suggest that balanced modality use does not guarantee superior performance, challenging the assumption that optimal multimodal models must use all modalities equally. Future work will systematically study, at scale, how audio information is leveraged, how it survives multimodal integration, and how task formulation (e.g., multiple choice vs. descriptive) influences its use.
7 Acknowledgments
This work is supported by the National Science Foundation under Grant Numbers 2504642 and 2504643. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.